Find the altitude of an isosceles trapezoid if its perimeter is 40cm, the area is 84cm^2 and each leg is 11cm long
Answer by advanced_Learner(501)

P=a+b+c+d
A=0.5(a+b)*h since a=b
A=0.5(2a)*h
P=(a+a)+c+d
40=2a+11+11
a=9
84=0.5(18)*h
h=28/3